# FlowMalTrans: Unsupervised Binary Code Translation for Malware Detection Using Flow-Adapter Architecture

### v1

@@ 1 Introduction
Malware is defined as software intended to damage computers or associated systems preda2008semantics. In recent years, many malware detection tools have been developed xie2024mmpalm, and their success largely depends on the underlying techniques. Signature-based detection identifies malware by matching patterns from known malware families sathyanarayan2008signature, but it often fails to detect altered or novel malware. In contrast, behavior-based detection analyzes program execution to identify suspicious behavior liu2011behavior, yet this method lacks scalability.
The use of deep learning methods like LSTM and CNNs in malware detection has garnered significant interest because of their strong performance sewak2018investigation; gopinath2023comprehensive. However, there are two major challenges that hinder the broader adoption of deep learning in this domain.
Challenges. Malware infects a broad range of Instruction Set Architectures (ISAs). With the rise in cyberattacks on IoT devices, malware has begun targeting an increasing number of ISAs davanian2022malnet; caviglione2020tight. Hence, it is essential to enable malware detection across ISAs. Currently, there are numerous ISAs available wang-security23. Due to this wide variety, training one model for each ISA demands substantial time and effort. In addition, deep learning models rely heavily on having enough malware samples during training, which is problematic for low-resource ISAs where such data is limited.
Our Approach. Malware is generally closed-source whose original source code is typically inaccessible. What is available is the binary form of the malware. Once disassembled, this binary can be represented in assembly language. Based on this observation, we propose leveraging Neural Machine Translation (NMT) techniques artetxe2017unsupervised to facilitate malware analysis.
@@
In order to capture ISA-specific syntax and semantic information, we leverage normalizing flows (NFs), a class of invertible transformations that models complex distributions while maintaining tractability. Traditional translation models often struggle with capturing complex dependencies and handling the variation in data distributions. NFs address this challenge by modeling complex, high-dimensional distributions through invertible transformations, enabling more efficient learning of intricate relationships between source and target data.
We design FlowMalTrans, which is a flow-adapter-based unsupervised binary code translation architecture. To bridge the gap between the latent space of the source and target ISA, FlowMalTrans learns an invertible transformation of binary representations from one ISA to another, facilitating robust feature alignment across different ISAs and enhancing translation fidelity. FlowMalTrans operates in a completely unsupervised manner, eliminating the need for parallel datasets. Notably, the training of FlowMalTrans does not require any malware samples and relies only on binaries compiled from the abundance of opensource programs. Despite never encountering any malware samples during training, FlowMalTrans is still capable of translating malware across ISAs with high translation quality.
Results. We have implemented FlowMalTrans, and evaluated its performance on seven ISAs, including X86-64, i386, ARM32, ARM64, MIPS32, PPC32 and M68K. FlowMalTrans achieves higher BLEU score than the baseline method UnsuperBinTrans ahmad-luo-2023-unsupervised. Furthermore, we apply FlowMalTrans to the malware detection task and achieve exceptional results, highlighting its superior translation quality, which enables highly effective malware detection across ISAs. Below we highlight our contributions:
  • We propose FlowMalTrans, a novel unsupervised approach to translate binaries across different ISAs. FlowMalTrans addresses the data scarcity issue by enabling the detection of malware in low-resource ISAs using a model trained on the high-resource ISA.
@@ 2.1 Malware Detection
Signature-Based Detection. Conventional malware detection methods primarily use signature-based approaches, working by recognizing malicious code patterns sebastio2020optimizing; rohith2021comprehensive; sathyanarayan2008signature; behal2010signature. However, these methods are ineffective against newly developed or altered malware that modifies its code to avoid being identified.
Behavior-Based Detection. This approach detects malware by examining program behavior during execution liu2011behavior; burguera2011crowdroid; aslan2021intelligent; saracino2016madam. However, it may produce false negatives when malicious actions are not activated during observation.
Deep Learning-Based Detection. Deep learning has been widely used for detecting malware. A range of models, including CNNs and LSTMs, have been utilized sewak2018investigation; gopinath2023comprehensive; wang2024learning. Nonetheless, these models generally demand substantial datasets of malware samples for training. Consequently, most existing methods concentrate on widely used ISAs like X86 and ARM, leaving low-resource ISAs relatively unaddressed.
@@ 2.2 Normalizing Flows
Normalizing flows (NFs) are a family of generative models that transform a simple base distribution into a complex target distribution via a series of invertible transformations ho2019flow++. The key idea is to apply a sequence of differentiable mappings to a latent variable. Different types of NFs include coupling layers dinh2014nice, autoregressive flows papamakarios2017masked, and continuous normalizing flows chen2018neural.
The flow adapter architecture has been explored in neural machine translation (NMT), such as variational NMT calixto2019latent; ma2019flowseq; eikema2019auto. Flow adapters leverage NFs to model complex posterior distributions in sequence-to-sequence models, allowing for more flexible latent representations and enhancing the robustness of translation shu2020latent.
@@ 2.3 Source Code Translation
Source code translation converts source code in one language to another. Early work uses transpilers or transcompilers andres2017transpiler; tripuramallu2024towards, which rely on handcrafted rules. However, they produce unidiomatic translations that prove hard for human programmers to read. Another issue is incomplete feature support, resulting in improper translations.
Deep learning have introduced new approaches to source code translation roziere2020unsupervised; weisz2021perfection; lachaux2020unsupervised. However, since malware is closed-source, source code translation is inapplicable for malware.
@@ 2.4 Program Analysis-Based Binary Code Translation
Several approaches leverage program analysis techniques to convert binary code from one ISA to another. They can be broadly categorized into static analysis-based and dynamic analysis-based. Static analysis-based translation analyzes and translates the binary code before execution shen2012llbt; cifuentes2000uqbt. But ensuring all paths are accurately translated is challenging.
Dynamic analysis-based translation performs translation during execution chernoff1998fx; ebcioglu2001dynamic, but it requires sophisticated runtime environments. Plus, both approaches face challenges related to architecture-specific features and encounter difficulties in achieving accuracy, performance, and compatibility.
Summary. We propose an unsupervised binary code translation approach leveraging deep learning techniques. The state of the art, UnsuperBinTrans ahmad-luo-2023-unsupervised, is the first and only existing work applying deep learning to binary code translation. However, it has several limitations, which our model FlowMalTrans aims to overcome (Section 3.2). Our evaluation demonstrates that FlowMalTrans outperforms UnsuperBinTrans in binary code translation and achieves superior malware detection performance.
@@ 3.2 Limitations of the State of the Art
UnsuperBinTrans ahmad-luo-2023-unsupervised is the first and only existing work that applies deep learning techniques to binary code translation. While InnerEye zuo2018neural applies NMT techniques for binary code similarity comparison, it does not perform binary code translation across ISAs. Instead, it employs two encoders from NMT models to generate embeddings for binary pairs. UnsuperBinTrans, however, has several limitations, which our model aims to overcome.
@@ 4.1 Instruction Normalization
  • (R1): We use IDA Pro IDA to disassemble binaries, which generates dummy names ida-dummy. We replace dummy names with their respective prefixes. For example, off_ and seg_ represent offset pointer value and segment address value. They are replaced with <OFF> and <SEG>.
@@
Figure 2: Model architecture of FlowMalTrans.
@@ 4.2 Normalizing Flows
The state-of-the-art, UnsuperBinTrans, assumes the latent representations of source and target basic blocks share a common space and learns ISA-agnostic latent representations. However, this can be overly restrictive, as different ISAs exhibit distinct syntactic and structural characteristics that are not always easily aligned in a shared latent space. As a result, UnsuperBinTrans may lose critical ISA-specific features necessary for accurate translation, leading to suboptimal performance.
Our model FlowMalTrans assumes the representations are different due to ISA-specific characteristics. FlowMalTrans uses a pair of NFs (i.e., source flow and target flow) to model the distributions of basic-block latent representations in the source and target ISA. During translation, a latent code transformation is performed to transform the source representation $z_{src}$ into the target representation $z_{tgt}$, which is fed to the decoder to generate a better basic block in the target ISA. Figure 2 shows how NFs are employed in our model.
Modeling Latent Representations by NFs. We denote $z_{src}$ and $z_{tgt}$ as the latent representation of the source block and target block, respectively. Through NFs, we transform the distributions of the source and target representations to the base distribution $\epsilon$ (e.g., standard normal distribution), which can be viewed as the “true” underlying semantic space, abstracting away from ISA specifics. We denote such mappings as $\textbf{G}_{(z_{src}\rightarrow\epsilon)}$ and $\textbf{G}_{(z_{tgt}\rightarrow\epsilon)}$.
Designing NFs. We use Real-valued Non-volume Preserving (RealNVP), a type of NFs designed for efficient density estimation and sampling dinh2022density, to build the source and target flow.
@@
Therefore, we can learn a transformation $\textbf{G}_{(z_{src}\rightarrow z_{tgt})}$ by learning the transformation $\textbf{G}_{(z_{src}\rightarrow\varepsilon)}$ and $\textbf{G}_{(\varepsilon\rightarrow z_{tgt})}$. We can follow a similar procedure and learn $\textbf{G}_{(z_{tgt}\rightarrow z_{src})}$.
@@ 4.3 Encoder & Decoder for Code Translation
For binary code translation, different ISAs exhibit significant syntactic and semantic differences. To capture and align these variations, we propose a separated-encoder design, where each ISA is assigned a dedicated encoder. The source encoder and decoder are responsible for encoding and generating basic blocks in the source ISA, while the target encoder and decoder handle basic blocks in the target ISA. FlowMalTrans employs a multi-layer bi-directional Transformer vaswani2017attention to construct the encoders and decoders for the source and target ISA, as shown in Figure 2.
Encoding. The source encoder processes the source basic block $\textbf{B}_{src}=\{\boldsymbol{b}_{0},\cdots,\boldsymbol{b}_{S}\}$, and generates the hidden representations $\{\boldsymbol{h}_{0},\cdots\boldsymbol{h}_{S}\}$. The basic block-level representation $z_{src}$ is computed using Equation (4). The target encoder follows a similar process to encode the target basic block.
@@
${\qquad\qquad\qquad+\text{mean-pool}([\boldsymbol{h}_{0},\dots,\boldsymbol{h}_{S}])+\boldsymbol{h}_{0})}$ (4)
@@
Decoding. To capture the semantics and mitigate improper alignments between the source and target ISA, we follow the procedure in setiawan2020variational to generate the decoded representation:
@@ 5.1 Experimental Settings
We implement FlowMalTrans using Transformer with $64$ hidden units, $4$ heads, ReLU activations, a dropout rate of $0.1$, and learned position embeddings. Appendix A.4 presents the details.
@@
  • Baseline Model 1: UnsuperBinTrans ahmad-luo-2023-unsupervised is the first and only existing work that focuses on binary code translation leveraging deep learning (Section 3.2).
  • Baseline Model 2: UniMap wang-security23 also aims to resolve the data scarcity issue in malware detection. UniMap learns transfer knowledge to enable the reuse of a model across ISAs. In contrast, we focus on binary code translation to enable model reuse.
  • Baseline Model 3: CrossIns2Vec wang2024learning, similar to UniMap, aims to learn transfer knowledge to enable model reuse.
  • Baseline Model 4: IR-based malware detection model. Intermediate representation (IR) can abstract away ISA differences and represent binaries in a uniform style vex-ir.
@@ 5.2 Training FlowMalTrans
Training Datasets. We collect various opensource programs, including openssl, binutils, findutils, and libgpg-error, which are widely used in prior binary analysis works asm2vec-Ding; li2021palmtree. We compile them for each ISA using GCC with different optimization levels (O0-O3) and disassemble the binaries using IDA Pro IDA to extract basic blocks, which are normalized and deduplicated. Finally, we create a training dataset for each ISA. The adequacy of these datasets is evaluated in Appendix A.3.
@@
  • The vocabulary size discrepancy between the source and target ISA should $15\%$. A large vocabulary discrepancy can lead to an imbalanced learning problem, resulting in inefficiencies and overfitting gowda2020finding.
  • The vocabulary size of each ISA should <12k<12k, to balance capturing word semantics with computational resource constraints. A large vocabulary size can negatively impact model performance due to increased complexity and sparse token distributions jean2014using.

### v2

@@ 1 Introduction
Malware is defined as software intended to damage computers or associated systems Preda et al. (2008). In recent years, many malware detection tools have been developed Xie et al. (2024), and their success largely depends on the underlying techniques. Signature-based detection identifies malware by matching patterns from known malware families Sathyanarayan et al. (2008), but it often fails to detect altered or novel malware. In contrast, behavior-based detection analyzes program execution to identify suspicious behavior Liu et al. (2011), yet this method lacks scalability.
The use of deep learning methods like LSTM and CNNs in malware detection has garnered significant interest because of their strong performance Sewak et al. (2018); Gopinath and Sethuraman (2023). However, there are two major challenges that hinder the broader adoption of deep learning in this domain.
Challenges. Malware infects a broad range of Instruction Set Architectures (ISAs). With the rise in cyberattacks on IoT devices Chen et al. (2025); Ma et al. (2023); Tsai et al. (2024), malware has begun targeting an increasing number of ISAs Davanian and Faloutsos (2022); Caviglione et al. (2020). Hence, it is essential to enable malware detection across ISAs. Currently, there are numerous ISAs available Wang et al. (2023a). Due to this wide variety, training one model for each ISA demands substantial time and effort. In addition, deep learning models rely heavily on having enough malware samples during training, which is problematic for low-resource ISAs where such data is limited.
Our Approach. Malware is generally closed-source whose original source code is typically inaccessible. What is available is the binary form of the malware. Once disassembled, this binary can be represented in assembly language. Based on this observation, we propose leveraging Neural Machine Translation (NMT) techniques Artetxe et al. (2018) to facilitate malware analysis.
@@
In order to capture ISA-specific syntax and semantic information, we leverage normalizing flows (NFs), a class of invertible transformations that models complex distributions while maintaining tractability. Traditional translation models often struggle with capturing complex dependencies and handling the variation in data distributions. NFs address this challenge by modeling complex, high-dimensional distributions through invertible transformations, enabling more efficient learning of intricate relationships between source and target data. By doing so, we ensure a smooth and structured mapping between ISAs, leading to more precise and reliable translations. Furthermore, normalizing flows enable efficient sampling and density estimation, which improves the adaptability of our model to unseen binaries and enhances generalization across different architectures.
We design FlowMalTrans, which is a flow-adapter-based unsupervised binary code translation architecture. This design allows for robust feature alignment between different ISAs, improving translation fidelity and ultimately enhancing malware detection across ISAs. To bridge the gap between the latent space of the source and target ISA, FlowMalTrans learns an invertible transformation of binary representations from one ISA to another, facilitating robust feature alignment across different ISAs and enhancing translation fidelity. FlowMalTrans operates in a completely unsupervised manner, eliminating the need for parallel datasets. Notably, the training of FlowMalTrans does not require any malware samples and relies only on binaries compiled from the abundance of opensource programs. Despite never encountering any malware samples during training, FlowMalTrans is still capable of translating malware across ISAs with high translation quality.
Results. We have implemented FlowMalTrans, and evaluated its performance on seven ISAs, including X86-64, i386, ARM32, ARM64, MIPS32, PPC32 and M68K. FlowMalTrans achieves higher BLEU score than the baseline method UnsuperBinTrans Ahmad and Luo (2023). Furthermore, we apply FlowMalTrans to the malware detection task and achieve exceptional results, highlighting its superior translation quality, which enables highly effective malware detection across ISAs. Below we highlight our contributions:
  • We propose FlowMalTrans, a novel unsupervised approach to translate binaries across different ISAs. FlowMalTrans addresses the data scarcity issue by enabling the detection of malware in low-resource ISAs using a model trained on the high-resource ISA. Importantly, the training of FlowMalTrans does not require any malware samples, yet it is still capable of translating malware across ISAs and achieves high translation quality.
@@ 2.1 Malware Detection
Signature-Based Detection. Conventional malware detection methods primarily use signature-based approaches, working by recognizing malicious code patterns Sebastio et al. (2020); Rohith and Kaur (2021); Sathyanarayan et al. (2008); Behal et al. (2010). However, these methods are ineffective against newly developed or altered malware that modifies its code to avoid being identified.
Behavior-Based Detection. This approach detects malware by examining program behavior during execution Liu et al. (2011); Burguera et al. (2011); Aslan et al. (2021); Saracino et al. (2016). However, it may produce false negatives when malicious actions are not activated during observation.
Deep Learning-Based Detection. Deep learning has been widely used for detecting malware. A range of models, including CNNs and LSTMs, have been utilized Sewak et al. (2018); Gopinath and Sethuraman (2023); Wang et al. (2024); He et al. (2023); Lei et al. (2022); Zhang et al. (2024); Zhao et al. (2025). Nonetheless, these models generally demand substantial datasets of malware samples for training. Consequently, most existing methods concentrate on widely used ISAs like X86 and ARM, leaving low-resource ISAs relatively unaddressed.
@@ 2.2 Normalizing Flows
Normalizing flows (NFs) are a family of generative models that transform a simple base distribution into a complex target distribution via a series of invertible transformations Ho et al. (2019). The key idea is to apply a sequence of differentiable mappings to a latent variable. Different types of NFs include coupling layers Dinh et al. (2014), autoregressive flows Papamakarios et al. (2017), and continuous normalizing flows Chen et al. (2018).
The flow adapter architecture has been explored in neural machine translation (NMT), such as variational NMT Calixto et al. (2019); Ma et al. (2019); Eikema and Aziz (2019). Flow adapters leverage NFs to model complex posterior distributions in sequence-to-sequence models, allowing for more flexible latent representations and enhancing the robustness of translation Shu et al. (2020).
@@ 2.3 Source Code Translation
Source code translation converts source code in one language to another. Early work uses transpilers or transcompilers Andrés and Pérez (2017); Tripuramallu et al. (2024), which rely on handcrafted rules. However, they produce unidiomatic translations that prove hard for human programmers to read. Another issue is incomplete feature support, resulting in improper translations.
Deep learning have introduced new approaches to source code translation Roziere et al. (2020); Weisz et al. (2021); Lachaux et al. (2020). However, since malware is closed-source, source code translation is inapplicable for malware.
@@ 2.4 Program Analysis-Based Binary Code Translation
Several approaches leverage program analysis techniques to convert binary code from one ISA to another. They can be broadly categorized into static analysis-based and dynamic analysis-based. Static analysis-based translation analyzes and translates the binary code before execution Shen et al. (2012); Cifuentes and Van Emmerik (2000). But ensuring all paths are accurately translated is challenging.
Dynamic analysis-based translation performs translation during execution Chernoff et al. (1998); Ebcioglu et al. (2001), but it requires sophisticated runtime environments. Plus, both approaches face challenges related to architecture-specific features and encounter difficulties in achieving accuracy, performance, and compatibility.
Summary. We propose an unsupervised binary code translation approach leveraging deep learning techniques. The state of the art, UnsuperBinTrans Ahmad and Luo (2023), is the first and only existing work applying deep learning to binary code translation. However, it has several limitations, which our model FlowMalTrans aims to overcome (Section 3.2). Our evaluation demonstrates that FlowMalTrans outperforms UnsuperBinTrans in binary code translation and achieves superior malware detection performance.
@@ 3.2 Limitations of the State of the Art
UnsuperBinTrans Ahmad and Luo (2023) is the first and only existing work that applies deep learning techniques to binary code translation. While InnerEye Zuo et al. (2018) applies NMT techniques for binary code similarity comparison, it does not perform binary code translation across ISAs. Instead, it employs two encoders from NMT models to generate embeddings for binary pairs. UnsuperBinTrans, however, has several limitations, which our model aims to overcome.
@@ 4.1 Instruction Normalization
  • (R1): We use IDA Pro IDA (2023) to disassemble binaries, which generates dummy names Dummy name (2023). We replace dummy names with their respective prefixes. For example, off_ and seg_ represent offset pointer value and segment address value. They are replaced with <OFF> and <SEG>.
@@
Figure 2: Model architecture of FlowMalTrans, which contains a pair of flow adapters for modeling the distributions of basic block representations in the source ISA and target ISA, respectively.
@@ 4.2 Normalizing Flows
We consider basic blocks as sentences, and instructions as words. The prior state-of-the-art, UnsuperBinTrans, assumes that the latent representations of source and target basic blocks share a common semantic space, and learns ISA-agnostic latent representations, which abstract away ISA-specific details. However, this can be overly restrictive, as different ISAs exhibit distinct syntactic and structural characteristics that are not always easily aligned in a shared latent space. By enforcing ISA-agnostic representations, UnsuperBinTrans may lose critical ISA-specific features necessary for accurate translation, leading to suboptimal performance.
Our model FlowMalTrans assumes the representations are different due to ISA-specific characteristics. Thus, they are modeled separately for each ISA, which makes it possible to better capture basic-block semantics in an ISA-specific manner. Specifically, FlowMalTrans uses a pair of NFs (i.e., source flow and target flow) to model the distributions of basic-block latent representations in the source and target ISA. During translation, a latent code transformation is performed to transform the source representation $z_{src}$ into the target representation $z_{tgt}$, which is fed to the decoder to generate a better basic block in the target ISA. Figure 2 shows how NFs are employed in our model.
Modeling Latent Representations by NFs. We use a pair of NFs to model the distribution of the basic block latent representations in different ISAs, i.e., $p_{z_{src}}(z_{src})$ and $p_{z_{tgt}}(z_{tgt})$, where $z_{src}$ and $z_{tgt}$ are the latent representation of the source basic block and target basic block, respectively. Through NFs, we transform the distributions of the source and target representations to the base distribution $\epsilon$ (e.g., standard normal distribution), which can be viewed as the “true” underlying semantic space, abstracting away from ISA specifics. We denote such mappings as $\textbf{G}_{(z_{src}\rightarrow\epsilon)}$ and $\textbf{G}_{(z_{tgt}\rightarrow\epsilon)}$.
Designing NFs. We use Real-valued Non-volume Preserving (RealNVP), a type of NFs designed for efficient density estimation and sampling Dinh et al. (2022), to build the source and target flow.
@@
Therefore, we can learn a transformation $\textbf{G}_{(z_{src}\rightarrow z_{tgt})}$ by learning the transformation $\textbf{G}_{(z_{src}\rightarrow\varepsilon)}$ and $\textbf{G}_{(\varepsilon\rightarrow z_{tgt})}$. We also notice that $\textbf{G}_{(z_{src}\rightarrow z_{tgt})}$ and $\textbf{G}_{(z_{tgt}\rightarrow z_{src})}$ are invertible because they are compositions of two invertible mappings. Thus, We can follow a similar procedure and learn $\textbf{G}_{(z_{tgt}\rightarrow z_{src})}$.
@@ 4.3 Encoder & Decoder for Code Translation
For binary code translation, different ISAs exhibit significant syntactic and semantic differences. To capture and align these variations, we propose a separated-encoder design, where each ISA is assigned a dedicated encoder. The source encoder and decoder are responsible for encoding and generating basic blocks in the source ISA, while the target encoder and decoder handle basic blocks in the target ISA. FlowMalTrans employs a multi-layer bi-directional Transformer Vaswani (2017) to construct the encoders and decoders for the source and target ISA, as shown in Figure 2.
Encoding. Here we describe how the source encoder encodes the source basic block and generates the source basic block representation $z_{src}$. A similar procedure is applied to the target basic block representation generation.
The source encoder processes the source basic block $\textbf{B}_{src}=\{\boldsymbol{b}_{0},\cdots,\boldsymbol{b}_{S}\}$, and generates the hidden representations $\{\boldsymbol{h}_{0},\cdots\boldsymbol{h}_{S}\}$. The basic block-level representation $z_{src}$ is computed using Equation (4). The target encoder follows a similar process to encode the target basic block.
@@
$+\text{mean-pool}([\boldsymbol{h}_{0},\dots,\boldsymbol{h}_{S}])$
$+\boldsymbol{h}_{0})$ (4)
@@
In contrast, the prior work UnsuperBinTrans indiscriminately uses the same representational space for both source and target, without performing latent code transformation.
Decoding. To capture the semantics and mitigate improper alignments between the source and target ISA, we follow the procedure in Setiawan et al. (2020) to generate the decoded representation:
@@ 5.1 Experimental Settings
We implement FlowMalTrans using Transformer with $64$ hidden units, $4$ heads, ReLU activations, a dropout rate of $0.1$, and learned position embeddings. Appendix A.4 presents the details. All the experiments were conducted on a computer with a $64$-bit 3.6 GHz Intel Core i9-CPU, a Nvidia GeForce RTX 4090 and 64GB RAM.
@@
  • Baseline Model 1: UnsuperBinTrans Ahmad and Luo (2023) is the first and only existing work that focuses on binary code translation leveraging deep learning (Section 3.2).
  • Baseline Model 2: UniMap Wang et al. (2023a) also aims to resolve the data scarcity issue in malware detection. UniMap learns transfer knowledge to enable the reuse of a model across ISAs. In contrast, we focus on binary code translation to enable model reuse.
  • Baseline Model 3: CrossIns2Vec Wang et al. (2024), similar to UniMap, aims to learn transfer knowledge to enable model reuse.
  • Baseline Model 4: IR-based malware detection model. Intermediate representation (IR) can abstract away ISA differences and represent binaries in a uniform style angr (2024).
@@ 5.2 Training FlowMalTrans
Training Datasets. We collect various opensource programs, including openssl, binutils, findutils, and libgpg-error, which are widely used in prior binary analysis works Ding et al. (2019); Li et al. (2021). We compile them for each ISA using GCC with different optimization levels (O0-O3) and disassemble the binaries using IDA Pro IDA (2023) to extract basic blocks, which are normalized and deduplicated. Finally, we create a training dataset for each ISA. The adequacy of these datasets is evaluated in Appendix A.3.
@@
  • The vocabulary size discrepancy between the source and target ISA should $15\%$. A large vocabulary discrepancy can lead to an imbalanced learning problem, resulting in inefficiencies and overfitting Gowda and May (2020).
  • The vocabulary size of each ISA should <12k<12k, to balance capturing word semantics with computational resource constraints. A large vocabulary size can negatively impact model performance due to increased complexity and sparse token distributions Jean et al. (2014).